OpenStudy (anonymous):

Consider The polnomial: P(x)=40x^3+116x^2-24x-36.....with one of its zeros being x1=-3. Find the remaining zeros by using synthetic division. answer choices: 1.x2,3=-1/2,4/5 2.None of these 3.x2,3=-1/4,4/5 4.x2,3=-1/2,3/5 5x2,3=-1/4,3/5 (can someone help me work this out work provided so i can understand this question more in case or future reference? plz and ty ^.^)

OpenStudy (dumbcow):

x = -3 --> x+3 = 0 So you have to divide P(x) by x+3 to solve for remaining 2 zeros using synthetic division, put -3 on outside and the coefficients on inside remember to multiply then add, multiply then add ... -3| 40 116 -24 -36 -120 12 36 --------------------- 40 -4 -12 0 The resulting numbers are the coefficients of the quotient, notice the remainder is zero P(x) = (x+3)(40x^2 -4x -12) = 0 Solve the quadratic 40x^2 -4x -12 = 0 Divide equation by 4 10x^2 -x -3 = 0 Factor or use quadratic formula 10x^2 +5x -6x -3 = 0 5x(2x +1) -3(2x +1)=0 (2x+1)(5x -3) = 0 2x+1 = 0 --> x = -1/2 5x-3=0 --> x = 3/5